OpenStudy (anonymous):

Can someone help me with some pH? Calculate the concentration of sulfuric acid in a solution with a pH of 2.0 Can you explain it please?

OpenStudy (anonymous):

you asked that few minutes ago

OpenStudy (anonymous):

Can you explain it? it is a little difficult when you last showed me

OpenStudy (anonymous):

right, i partly understood it because my teacher wrote what you said, but its confusing me a bit

OpenStudy (anonymous):

she wrote [h+] = 10-pH=10^2-0.01MH+1mol H2So4/2mol H+

OpenStudy (anonymous):

ok so pH is a relative measure (logaritmic scale) of acidity/basicity of solution, so we measure it by determining concentration of H+ ions, as H2SO4 dissolves completely it releases 2 H+ ions you follow me?

OpenStudy (anonymous):

if it is a strong acid, it dissolves completely right

OpenStudy (anonymous):

thats why it dissolves completely releasing the 2 H'+?

OpenStudy (anonymous):

right, so you have pH which tells you concentration of those H+ ions, right? so you calculate concentration of those H+ ions from this equation: pH = - log [H+] --> antilog pH = [H+] where antilog is 10 ^ -pH so now you know concentration og H+ ions is 0,01 right?

OpenStudy (anonymous):

and as H2SO4 has 2 H+ ions you need to divide that concentration in half to get concentration of H2SO4, do you get it now?

OpenStudy (anonymous):

Oh, so is that where you get the 10^-2, which is .01. yes i see

OpenStudy (anonymous):

can you explain why you need to divide by half?

OpenStudy (anonymous):

cause H2SO4 dissociates and gives 2 H+ ions so ratio is 1:2 get it?

OpenStudy (anonymous):

so you just want to know for 1?

OpenStudy (anonymous):

yes cause you are asked to calculate concentration of H2SO4 right? and 2 H+ ions go in to forme H2SO4...